OpenStudy (anonymous):

Express the given repeating decimal .7 as a quotiant of integers and reduce if possible?

OpenStudy (anonymous):

What if it isn't possible? Does there exist a and b such that a/b = 0.77777777777...?

OpenStudy (anonymous):

integers a and b*

OpenStudy (anonymous):

well according to my book it said to let a given number be n and to multiply each side by 100 ending with 100n = 70.7 then subtract one n and .7 and you have 99n=70/99 but it doesn't seem right

OpenStudy (across):

I found one, but it wasn't easy. :( \[\frac{7}{9}\]

OpenStudy (across):

I wish I could explain it how I got there...

OpenStudy (anonymous):

Brute force?

OpenStudy (across):

No, I started analyzing repeating decimals which I already knew, such as 10/3 and played around with them to get a repeating 7.

OpenStudy (anonymous):

well. if you multiply 0.77777777... 9 times you get a whole number

hero (hero):

There is a method to these. I just don't remember

OpenStudy (anonymous):

Let x=.77777777... Then 10x=7.777777.... But 7.777777...= 7+.777777... = 7 + x So we have 7 + x = 10x which gives 7 = 9x which gives x= 7/9.

OpenStudy (anonymous):

In general (just as a short cut), if you have a repeating decimal, say: 0.abcabcabcabc.... Then the fraction will be abc/999 Its one complete cycle over the same number of 9s. Examples: .4545454545... = 45/99= 5/11 .1234123412341234... = 1234/9999

hero (hero):

You should get 10 or more medals for that outstanding explanation